# https://github.com/beichenzbc/Long-CLIP

### v2

@@ 3.2 Knowledge Preserving Stretching
In this specific task, the straightforward strategy of linear interpolation may not be the most suitable approach. This is because the majority of training texts are likely to be considerably shorter than 77 in the original CLIP model. As a result, the lower positions have been well-trained and can effectively indicate absolute positions. On the other hand, the higher positions have not received sufficient training and can only provide a rough estimation of relative positions. Consequently, the cost of interpolating the lower positions is much higher compared to interpolating the higher ones, as doing so is more likely to disrupt the well-established representation of absolute position.
@@ 3.3 Fine-tuning with Primary Component matching
However, simply fine-tuning the model with long captions will lead to a degradation on the short-text capability, especially in classification task. This means our model still can’t handle both long caption and short caption flexibly. The underlying reason is that a capable model should not only capture various attributes but also understand their relative relationships and differing importance. Merely fine-tuning the model with long captions may push it to another extreme, where it attempts to encompass all attributes in a single image without distinguishing their respective importance.
To this end, we propose a Primary Component matching strategy in long-text fine-tuning to both unlock the long-text capability and maintain the short-text capability. Apart from aligning the fine-grained feature of an image with its corresponding long caption, we extract a coarse-grained image feature that focuses on capturing key attributes. This coarse-grained feature is then aligned with the corresponding short summary caption. By doing so, we require the model not only to capture detailed attributes but also to discern and prioritize the importance of different attributes.
@@
Finally, we apply our component-reconstruction function $\mathcal{F}^{-1}$ to reconstruct the image feature with only the key components after filtration and their corresponding importance as Eq. 5.
@@ 4.1 Experiment Setting
2) short-caption image-text retrieval. For traditional short-caption image-text retrieval, we use COCO2017 [14] dataset and Flickr30k [32] dataset. For COCO2017, we use the 5k validation set. For Flickr30k, to increase difficulty and avoid benchmark saturation, we use the whole 30k dataset instead of the 1k test set. Moreover, we also use a dataset collected from DeepFashion [15] which contains 200 images and 200 similar texts to evaluate the fine-grained ability.
@@
Training setting.  We use the ShareGPT4V [2] dataset as training dataset, which contains about 1M (long caption, image) pairs. The random 1k data is separated as an evaluation dataset. We fine-tune the CLIP model for 1 epoch with batch size 2048. Detailed hyper-parameter settings will be listed in supplementary material.
@@ 4.2 New Evaluation Dataset on Long Caption Text-Image Retrieval
Specifically, we manually select 200 similar images from Visual Genome dataset [10], all showcasing a busy urban view. Then, we use GPT-4V [18] to generate a long and complete sentence to describe the image, including types, colors and relative locations of the attributes. Finally, we manually check and correct the generated captions. The model can successfully match the images with the correct captions only if it successfully understands and models the detailed attributes in both modalities.
@@
After the first submission, we further scaled up Urban-200 into Urban-1k. The dataset has been released at https://huggingface.co/datasets/BeichenZhang/Urban1k. Urban-200 is used in the main paper. Detailed results about Urban-1k is shown in Sec. 6 in supplementary materials.
@@ 4.3 Comparing with CLIP Model
We compare our model with CLIP in the above three different downstream tasks. The interpolation ratio $\lambda_{2}$ in knowledge-stretching is set 4, extending the maximum input length to 248.
To evaluate the scalability of our method, we also fine-tune CLIP VIT-L/14 model using our knowledge-preserved stretching of positional embedding and primary component matching strategy as well.
Moreover, we also compare with a direct fine-tuning method discussed in Sec. 1, which simply interpolates each position with a fixed ratio $\lambda_{1}=3$ and only apply (long text, image) pairs in training. The training dataset, total training epochs and the hyper-parameter settings are the same.
The detailed result is shown in Tab. 3, Tab. 3, and Tab. 3, which shows that our method outperforms our baseline in most aspects. Specifically, for more fine-grained task, i.e. short or long text-image retrieval, both image and text feature extracted have to represent more detailed information of multiple attributes to distinguish from other similar candidates. Therefore, our fine-tuned model can significantly improve performance. For coarse-grained classification tasks, our method does not suffer from significant performance degradation like the simple approach thanks to our interpolation and training strategies, which help keep its short-text capability. This will be further discussed in Sec. 4.4
@@ 4.4 Ablation Study
Effectiveness of our core component. There are two core component strategy in our fine-tuning method. The first is the knowledge-preserved stretching of positional embedding, which prevents a distortion on well-established positional representation. The second is the primary component matching, aligning both coarse-grained and fine-grained information for a text-image pair. To demonstrate that both strategies can effectively improve the performance of the model, we tested the model’s accuracy on different tasks both with and without using these two strategies. For the knowledge-preserved stretching (KPS), we compare it with interpolating each position with a fixed ratio 3. For the primary component matching (PCM), we compare it with simply fine-tuning with (long caption, image) pairs. The result is demonstrated in Tab. 4. As can be seen, removing any of the strategies we proposed will result in a significant loss of short text capabilities.
Table 4: A comparison of whether to use the above two strategies or not. Best result is in bold.
@@
Different strategy keeping short-text capability. We leverage primary component matching during long-text fine-tuning to extract the fine-grained image feature and align both coarse-grained and fine-grained information of both modalities. And obviously, there exist multiple strategies to help maintain the short-text capability. To further validate the superiority of our method, we design and implement three different strategies.
@@ 4.5 Plug-and-Play in Image Generation using CLIP
CLIP is widely used in many downstream task, including segmentation [12, 31], detection [13, 7] and most commonly, in text-to-image generation models like Stable Diffusion [23] and DALLE-2 [21]. The pre-trained text encoder is often used to extract the text feature of an input to serve as a guidance during image generating. However, just like retrieval tasks, the image generation models using the original CLIP text encoder are still limited by the restriction on 77 tokens and the lack of long-text capability.
Indeed, thanks to the interpolation and matching strategy employed in our Long-CLIP model, we are able to maintain the latent space of the original CLIP model while enhancing its ability to represent fine-grained details in long texts. As a result, our Long-CLIP model exhibits a plug-and-play effect when integrated into existing models.
For example, we can replace the original CLIP text encoder with our Long-CLIP text encoder in Stable-Diffusion-V1-5 [23] with no additional training and keeping other components unchanged. The provided figure (Fig. 6) offers a visual demonstration of the examples, showcasing the enhanced long-text capability of our model. Specifically, for shorter and simpler text prompts, the images generated by our model closely resemble those generated by the original model when all other settings are identical. This demonstrates that the short-text capability of the original model is preserved in our approach. However, when presented with longer and more detailed text prompts, our model surpasses the original model by capturing previously overlooked details and including information that exceeds the original 77-token limit.
@@ 5 Conclusion
In this work, we have propose Long-CLIP, a strong and flexible CLIP model with long-text capability. Our Long-CLIP model can support text inputs of up to 248 tokens, and can better capture the detailed attributes and their interrelation, thus obtaining a large improvement on retrieval task. Moreover, our model keeps the performance on zero-shot classification and can replace the pre-trained CLIP encoder in a plug-and-play manner in image generation task.
@@
Scaling-up Potential. Due to the scarcity of long text-image pairs, we only leverage 1M (image, long-text) pairs in ShareGPT4V [2]. Although the amount of data used is relatively small, the use of long text has shown significant improvements. This gives us imagination on the scaling-up potential when a large amount of data used. As the sufficient long texts can provide complex information, covering world knowledge, object properties, spatial relationships, and even aesthetic evaluations, the upper limit of the model’s ability will be significantly improved.
@@
Beichen ZhangPan ZhangXiaoyi Dong
Yuhang ZangJiaqi WangCorresponding author.  § Work done during an internship in Shanghai AI Laboratory. This work is partially supported by the National Key R&D Program of China (2022ZD0160201), and Shanghai Artificial lntelligence Laboratory.
@@ 7 Insufficient Long-text Ability for CLIP-based Models
8 More Examples on Retrieval and Image Generation

### v3

@@ 3.2 Knowledge Preserving Stretching
In this specific task, the straightforward strategy of linear interpolation may not be the most suitable approach. This is because the majority of training texts are likely to be considerably shorter than 77 in the original CLIP model. Therefore the lower positions have been well-trained and can effectively indicate absolute positions, while the higher positions have not received sufficient training and can only provide a rough estimation of relative positions. Consequently, the cost of interpolating the lower positions is much higher compared to interpolating the higher ones, as doing so is more likely to disrupt the well-established representation of absolute position.
@@ 3.3 Fine-tuning with Primary Component matching
However, simply fine-tuning the model with long captions will lead to a degradation on the short-text capability. The underlying reason is that a capable model should not only capture various attributes but also understand their relative relationships and differing importance. Merely fine-tuning the model with long captions may push it to another extreme, where it attempts to encompass all attributes in a single image without distinguishing their respective importance.
To this end, we propose a Primary Component matching strategy in long-text fine-tuning to both unlock the long-text capability and maintain the short-text capability. Apart from aligning the fine-grained feature of an image with its long caption, we extract a coarse-grained image feature that focuses on capturing key attributes. This coarse-grained feature is then aligned with a short summary caption. By doing so, we require the model not only to capture detailed attributes but also to discern and prioritize the importance of different attributes.
@@
Finally, we apply our component-reconstruction function $\mathcal{F}^{-1}$ to reconstruct the image feature with only the key components and their importance as Eq. 5.
@@ 4.1 Experiment Setting
2) short-caption image-text retrieval. For traditional short-caption image-text retrieval, we use COCO2017 [15] and Flickr30k [33]. For COCO2017, we use the 5k validation set. For Flickr30k, to increase difficulty and avoid benchmark saturation, we use the whole 30k dataset instead of the 1k test set.
@@
Training setting.  We use the ShareGPT4V [2] dataset as training dataset, which contains about 1M (long caption, image) pairs. The random 1k data is separated as an evaluation dataset. We fine-tune for 1 epoch with batch size 2048. Detailed hyper-parameter settings will be listed in supplementary material.
@@ 4.2 New Evaluation Dataset on Long Caption Text-Image Retrieval
Specifically, we manually select 200 similar images from Visual Genome dataset [10], all showcasing a busy urban view. Then, we use GPT-4V [18] to generate a long and complete sentence to describe the image, including types, colors and relative locations of the attributes. The model can successfully match the images with the correct captions only if it successfully understands and models the detailed attributes in both modalities.
@@
After the first submission, we further scaled up Urban-200 into Urban-1k. The dataset has been released at https://huggingface.co/datasets/BeichenZhang/Urban1k. Urban-200 is used in the main paper. Detailed results about Urban-1k is shown in supplementary materials.
@@ 4.3 Comparing with CLIP Model
We compare our model with CLIP in the above three different downstream tasks. The interpolation ratio $\lambda_{2}$ in knowledge-stretching is set 4, extending the maximum input length to 248. To evaluate the scalability of our method, we also fine-tune CLIP ViT-L/14 model using our knowledge-preserved stretching of positional embedding and primary component matching strategy as well.
Moreover, we also compare with a direct fine-tuning method discussed in Sec. 1, which simply interpolates each position with a fixed ratio $\lambda_{1}=3$ and only apply (long text, image) pairs in training.
The detailed result is shown in Tab. 3, Tab. 3, and Tab. 3, which shows that our method outperforms our baseline in most aspects. Specifically, for fine-grained task, i.e. short or long text-image retrieval, both image and text feature extracted can represent more detailed information of multiple attributes to distinguish from other similar candidates. Therefore, our model can significantly improve performance. For coarse-grained classification tasks, our method doesn’t suffer from significant performance degradation like the simple approach thanks to our interpolation and training strategies. This will be discussed in Sec. 4.4
@@ 4.4 Ablation Study
Effectiveness of our core component. There are two core component strategy in our method. The first is the knowledge-preserved stretching of positional embedding, which prevents a distortion on well-established positional representation. The second is the primary component matching, aligning both coarse-grained and fine-grained information for a text-image pair. To demonstrate that both strategies can improve the performance of the model, we tested the model’s accuracy on different tasks both with and without using these two strategies. For the knowledge-preserved stretching (KPS), we compare it with interpolating each position with a fixed ratio 3. For the primary component matching (PCM), we compare it with simply fine-tuning with (long caption, image) pairs. The result is demonstrated in Tab. 4. As can be seen, removing any of the strategies we proposed will result in a significant loss of short text capabilities.
Table 4: A comparison of whether to use our two strategies. Best result is in bold.
@@
Different strategy keeping short-text capability. We leverage primary component matching during long-text fine-tuning. And obviously, there exist multiple strategies to help maintain the short-text capability. To further validate the superiority of our method, we design and implement three different strategies.
@@ 4.5 Plug-and-Play in Image Generation using CLIP
CLIP is widely used in many downstream task, including segmentation [12, 32], detection [13, 7] and text-to-image generation models like Stable Diffusion [24]. The pre-trained text encoder is often used to extract the text feature of an input to serve as a guidance during image generating. However, just like retrieval tasks, the image generation models using the original CLIP text encoder are still limited by the restriction on 77 tokens and the lack of long-text capability.
Indeed, thanks to the interpolation and matching strategy, we are able to maintain the latent space of the original CLIP model while enhancing its ability to represent fine-grained details in long texts. As a result, our Long-CLIP model exhibits a plug-and-play effect when integrated into existing models.
For example, we can replace the original CLIP text encoder with our Long-CLIP text encoder in Stable-Diffusion-V1-5 [24] with no additional training and keeping other components unchanged. The provided figure (Fig. 6) offers a visual demonstration of the examples, showcasing the enhanced long-text capability of our model. For short and simple text prompts, the images generated by our model closely resemble those generated by the original model. This demonstrates that the short-text capability of the original model is preserved in our approach. However, when presented with longer and more detailed text prompts, our model surpasses the original model by capturing previously overlooked details and including information that exceeds the original 77-token limit.
@@
After the first submission, we further use our Long-CLIP model in SDXL [20], which will be discussed in detail in supplementary materials.
@@ 5 Conclusion
We have propose Long-CLIP, a strong and flexible CLIP model with long-text capability. Our model can support text inputs of up to 248 tokens, and can better capture the detailed attributes, obtaining a large improvement on retrieval task. Moreover, our model keeps the performance on zero-shot classification and can replace the CLIP encoder in a plug-and-play manner in image generation task.
@@
Scaling-up Potential. Due to the scarcity of long text-image pairs, we only leverage 1M (image, long-text) pairs in ShareGPT4V [2]. This gives us imagination on the scaling-up potential when a large amount of data used. As the sufficient long texts can provide complex information, covering world knowledge, object properties, spatial relationships, and even aesthetic evaluations, the model’s ability may be significantly improved.
Acknowledgments
This work is partially supported by the National Key R&D Program of China (2022ZD0160201), and Shanghai Artificial Intelligence Laboratory.
@@ Acknowledgments
Beichen ZhangPan ZhangXiaoyi DongCorresponding author.  § Work done during an internship in Shanghai AI Laboratory.
Yuhang ZangJiaqi Wangfootnotemark:
@@ 7 Insufficient Long-text Ability for CLIP-based Models
8 Long-CLIP with SDXL
After the first submission, we further use our Long-CLIP model in Stable-Diffusion-XL [20] in a plug-and-play manner. Specifically, we replace the CLIP-L text encoder with our Long-CLIP-L, and only apply knowledge-preserved stretching (KPS) on Open-CLIP bigG text encoder due to heavy training cost.
The results are shown in Fig. 7, which shows our model can help SDXL break the 77 token limit with little reduction in the image quality.
Figure 7: Our Long-CLIP model can help SDXL break through the 77-token limit (marked in red \\) and capture the originally truncated attribute (marked in orange) with little reduction in the image quality.
9 Generalizability of Proposed Strategy
We apply our strategy, namely knowledge-preserved stretching and primary component matching, to fine-tune DeCLIP ViT-B/32 [14], the result also shows a consistent improvement.
Table 8: The performance of DeCLIP model and Long-DeCLIP model after fine-tuning. ‘Avg Cls’ means the average accuracy among 5 classification datasets in the main paper. Retrieval tasks are reported in ‘T2I/I2T’ format with R@5 of short-caption (COCO/Flickr) and R@1 of long-caption (Urban/ShareGPT4V).
Model  Avg Cls  COCO  Flickr  Urban  ShareGPT4V
DeCLIP 63.6  60.5/45.8  36.2/25.3  28.0/25.5  63.1/60.7
 Long-DeCLIP 63.9  61.1/50.7  36.8/33.8  54.0/51.0  84.1/84.3